OpenStudy (anonymous):

How do you divide 3/7¸2

OpenStudy (anonymous):

Or 3/7 divided by two

OpenStudy (anonymous):

3/7/2 = (3/7)/2 = (3/7)*(1/2) = 3/14

OpenStudy (anonymous):

Thank you so much!!